How many cups of pure water should be mixed with 1.5 cups of a 80​% acid solution to dilute it into a 22​% acid​ solution?
? cups ​(Type an integer or a​ decimal.)
A rectangular solar panel has a length that is 11 inches shorter than 3 times its width. If the perimeter of the panel is 178 ​inches, what are the dimensions of the​ panel?
The width of the panel is
? inches.
​(Simplify your​ answer.)

Answer by ankor@dixie-net.com(22740)

How many cups of pure water should be mixed with 1.5 cups of a 80​% acid solution to dilute it into a 22​% acid​ solution?
:
let w = amt of pure water
.8(1.5) = .22(w+1.5)
1.2 = .22w + .33
1.2 - .33 = .22w
.87 = .22w
w = .87/.22
w = 3.95 cups of water required
:
;
A rectangular solar panel has a length that is 11 inches shorter than 3 times its width.
L = 3W - 11
If the perimeter of the panel is 178 ​inches, what are the dimensions of the​ panel?
2L + 2W = 178
simplify, divide by 2
L + W = 89
replace L with (3W-11), (equation from the first statement)
(3W-11) + W = 89
3W + W = 89 + 11
4W = 100
W = 100/4
W = 25 inches is the width
then
3(25) - 11 = 64 inches is the width